OpenStudy (anonymous):

3600 cm/hr^2 = ? m/s^2

OpenStudy (anonymous):

=3600/(100x3600x3600) =1/360000 = 2.778x10^-6 m/s^2

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Ok I can explain.. you know, 1m=100cm so it is divided by 100 to convert it from "cm" to "m" next, you know in 1 hour there are 60 min and in 1 min there are 60 sec so 1 hour= 60x60 sec = 3600sec there is square of hours so i used 3600x3600 in denominator ..

OpenStudy (anonymous):

Thank you so much!